How many grams of water can be heated from 20.0C to 75C using 12500.0 Joules?
Answer #1

We know the relationship

Q = m * C * ∆T -----(1)

Q = heat , m = mass of substance, C = heat capacity

∆T = Temperature change = (75 - 20) oC = 55 oC

For liquid water, C = 4.184 J/g.oC

From(1), m = Q / C * ∆T = 12500 J / 4.184 J/g.oC * 55 oC = 12500 J / 230.12 J/g = 54.32 g

Mass of water can be heated = 54.32 g